3. If the sum of three numbers in an A.P. is 9 and their product is 24, then numbers are
(a) 2, 4, 6
(b) 1, 5, 3
(c) 2, 8, 4
(d) 2, 3, 4

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: d Explaination:Reason: Let three numbers be a – d, a, a + d ∴ a – d +a + a + d = 9 ⇒ 3a = 9 ⇒ a = 3 Also (a – d) . a . (a + d) = 24 ⇒ (3 -d) .3(3 + d) = 24 ⇒ 9 – d² = 8 ⇒ d² = 9 – 8 = 1 ∴ d = ± 1 Hence numbers are 2, 3, 4 or 4, 3, 2